OpenStudy (anonymous):

what is the value of the range of the function f(x)=x^2+2 for the domain value 1/4

OpenStudy (anonymous):

So the range value is whatever y= and the domain value is given as x=0.25. Right? f(x) and y are the same thing by the way. f(0.25)=0.25^2+2 y=0.25^2+2 y=0.5+2 y=2.5